OpenStudy (anonymous):

Two cars leave town going in opposite directions. One car is traveling 70 miles per hour (mph), and the other car is traveling 50 mph. How long will it take before the cars are 240 miles apart?

OpenStudy (anonymous):

Each hour, one moves 50 miles in one direction, and the other moves 70 miles the opposite way. The total separation after one hour is 70 + 50 = 120 miles. In general, the total separation will be 120 miles * (# of hours). So how many hours will it take to get 240 miles of separation?

OpenStudy (anonymous):

would it be 2 hours since you need to get to 240 miles?

OpenStudy (anonymous):



OpenStudy (anonymous):

right, they move apart 120 miles each hour, so the equation is d = 120 * t where t is the number of hours and d is distance apart the problem asks for t when d = 240 240 = 120 * t 240/120 = t t = 2 hours Good work :)

OpenStudy (anonymous):

thx! @JakeV8

OpenStudy (anonymous):

Glad to help... is it making sense? You seem pretty good at these :)

OpenStudy (anonymous):

yea its making sense. could you help me at another one?

OpenStudy (anonymous):

@JakeV8

OpenStudy (anonymous):

sure... maybe post it as a new question?